. Assume in a market individual firm supply function of a commodity X is given by Sx=-10 + 4Px and individual demand function is Dx =6 -2Px. There are 10 suppliers of commodity X with identical supply function and there are 100 buyers of the commodity X with identical demand function. Find market equilibrium price and quantity demanded.


1
Expert's answer
2022-01-17T12:40:03-0500

Solution:

Market demand function = 100(6 – 2Px) = 600 – 200Px

Market supply function = 10(-10 + 4Px) = -100 + 40Px

 

At equilibrium: Market demand = market supply

600 – 200Px = -100 + 40Px

600 + 100 = 40Px + 200Px

700 = 240Px

Px = 2.92

Market equilibrium price = 2.92

 

Substitute to derive market equilibrium quantity demanded:

Market demand = 600 – 200Px = 600 – 200(2.92) = 600 – 584 = 16

Market equilibrium quantity demanded = 16